OpenStudy (anonymous):

Calculus: If x > 0, show that there exists a n in the naturals such that 1/(2^n) < n. My work will be included in the following comment.

OpenStudy (anonymous):

Here is my attempt at a proof: Proof by contradiction: Let \[x \in \mathbb{R} \mbox{ s.t } x > 0\]. Assume there is not a \[n \in \mathbb{N} \mathbb{ s.t. } \frac{1}{2^{n}} < x\]. That is assume \[\forall n \in n, \frac{1}{2^{n}} \geq x \mbox{ for } x > 0\] Pick any \[x \in \mathbb{R} > 0\] Then, \[\frac{1}{2^{n}} \geq x \] for some n. This is a contradiction. \[\therefore \exists n \in \mathbb{N} \mbox{ s.t. } \frac{1}{2^{n}} < x \] \[\square\] Is this looking accurate?

OpenStudy (anonymous):

Do we need to Prove this: 1/2^n <x for x>0 and n>0

OpenStudy (anonymous):

use whatever that axiom is called, the one that says there is no largest integer

OpenStudy (anonymous):

for all \(x\) there is an \(n\) such chat \(2^n>\frac{1}{x}\)